Assume that it is now January 1, 2017. Wayne-Martin Electric Inc. (WME) has developed a solar panel capable of generating 200% more electricity than any other solar panel currently on the market. As a result, WME is expected to experience a 14% annual growth rate for the next 5 years. Other firms will have developed comparable technology by the end of 5 years, and WME's growth rate will slow to 4% per year indefinitely. Stockholders require a return of 12% on WME's stock. The most recent annual dividend (D0), which was paid yesterday, was $1.88 per share.

  1. Calculate the value of the stock today, Proceed by finding the present value of the dividends expected at the end of 2017, 2018, 2019, 2020, and 2021 plus the present value of the stock price that should exist at the end of 2021. The year-end 2021 stock price can be found by using the constant growth equation. Notice that to find December 31, 2021, price, you must use the dividend expected in 2022, which is 4% greater than the 2021 dividend.
    Round your answer to the nearest cent. Do not round your intermediate calculations.
    $ ?
  2. Calculate the expected dividend yield (D1/P0), capital gains yield, and total return (dividend yield plus capital gains yield) expected for 2017. (Assume that    and recognize that the capital gains yield is equal to the total return minus the dividend yield.) Round your answers to two decimal places. Do not round your intermediate calculations.

    D1/P0 =  %

    Capital gains yield =  %

    Expected total return =  %

Please Help!

Thank you!
Answer #1

a. D0=Current Dividend =$1.88

g= Dividend growth rate next 5 years =14%=0.14

Expected dividend in Year 1(Next year)=D1=D0*(1+g)=1.88*1.14=$2.14

Expected dividend in Year 2=D2=D1*(1+g)=2.14*1.14=$2.44

Expected dividend in Year 3=D3=D2*(1+g)=2.44*1.14=$2.79

Expected dividend in Year 4=D4=D3*(1+g)=2.79*1.14=$3.18

Expected dividend in Year 5=D5=D4*(1+g)=3.18*1.14=$3.62

Growth rate after year 5=g1=4%=0.04

Expected dividend in Year 6=D6=D5*(1+g1)=3.62*1.04=$3.76

R=Required return=12%=0.12

Price in Year 5 =P5=D6/(R-g1)

P5=3.76/(0.12-0.04)=$47.06

Value of the stock today =Present Value of future cash flows=

Present Value of cash flow=(Cash Flow)/((1+R)^N)

R=Required Return=0.12, N=year of cash flow

Year(N) Cash Flow(CF) Present Value(PV=CF/(1.12^N)
1 D1 $2.14 $1.91
2 D2 $2.44 $1.95
3 D3 $2.79 $1.98
4 D4 $3.18 $2.02
5 D5 $3.62 $2.05
5 P5 $47.06 $26.70
SUM $36.62

Value of stock today=$36.62

b. Dividend yield for 2017=Dividend/current price =1.91/36.62=0.0523=5.23%

Capital Gain Yield in 2017=12%-5.23%=6.77%

Expected Total Return =5.23%+6.77%=12%